ADNAN MUHAMMED

Grade 12,

What is the SI unit of planks contantand and find dimensions of planks constant, Energy, and force

What is the SI unit of planks contantand and find dimensions of planks constant, Energy, and force

Grade:11

3 Answers

Arun
4 years ago
Dear student
 
E = h* v
 
h is planck constant 
 
dimension of h = ML^2 T^-1
 
hence unit is Joule- sec
 
Hope it helps
 
Vikas TU
4 years ago
Dear student 
Planck's constant, symbolized h, relates the energy in one quantum (photon) of electromagnetic radiation to the frequency of that radiation. In the International System of units (SI), the constant is equal to approximately 6.626176 x 10-34 joule-seconds.
Its  relates the energy carried by a photon to its frequency. A photon's energy is equal to its frequency multiplied by the Planck constant.
E = h* v
Dimention of Planks constant = h = ML^2 T^-1
Hope this helps 
Good luck 
Cheers 
 
RISHIKA
3 years ago
HEY   FRIEND
FOR FINDING PLANKS CONSTANT FIRST WE HAD TO KNOW ITS FORMULA
SO IT IS AS                E = h*v                                 h=E/v
NOW,    WE KNOW THAT (V)IS FREQUENCY WHICH IS EQUAL TO THE “ RECIPROCAL OF TIME”
THEREFORE,    h=E*T
                          (WE KNOW THAT SI UNIT OF E=JOULE &  T =SECOND)
THUS,h=JOULE  SECOND
 DIMENSION= ML^2T^-1
I THOUGHT IT IS HELPFUL
BE HAPPY